# A Tie-breaking based Local Search Algorithm for Stable Matching Problems

### v1

@@ Abstract
The stable marriage problem with incomplete lists and ties (SMTI) and the hospitals/residents problem with ties (HRT) are important in matching theory with broad practical applications. In this paper, we introduce a tie-breaking based local search algorithm (TBLS) designed to achieve a weakly stable matching of maximum size for both the SMTI and HRT problems. TBLS begins by arbitrarily resolving all ties and iteratively refines the tie-breaking strategy by adjusting the relative order within ties based on preference ranks and the current stable matching. Additionally, we introduce TBLS-E, an equity-focused variant of TBLS, specifically designed for the SMTI problem. This variant maintains the objective of maximizing matching size, while enhancing equity through two simple modifications. In comparison with ten other approximation and local search algorithms, TBLS achieves the highest matching size, while TBLS-E exhibits the lowest sex equality cost. Significantly, TBLS-E preserves a matching size comparable to that of TBLS. Both our algorithms demonstrate faster computational speed than other local search algorithms in solving large-sized instances.
@@ 1 Introduction
A natural variant of SM and HR incorporates ties and incomplete preference lists, allowing the agents to express indifference and exclude unacceptable members. These variants are known as the stable marriage problem with incomplete lists and ties (SMTI) [19] and the hospitals/residents problem with ties (HRT) [17]. Recently, the SMTI and HRT problems have received significant attention from researchers due to their relevance in diverse applications such as the child-adoption matching problem [7, 31] and the student-project allocation problem [1, 2, 25]. When ties are present in the preference lists, three criteria are used to define the stability of a match: weak stability, strong stability, and super-stability [15, 29]. Among these definitions, weak stability has been the focus of extensive research [4, 7, 13, 11].
Beyond achieving stability, objectives also include maximizing matching size and ensuring fairness. For a given SM or HR problem, all stable matchings have the same size [10]. A stable matching can be easily found using the Gale-Shapley algorithm (GS) proposed by [9]. For the SMTI or HRT problem, a weakly stable matching can be obtained by applying GS after breaking all ties. However, different methods of breaking ties generally result in stable matchings of varying sizes [17, 18]. Consequently, a natural objective is to find a maximum cardinality weakly stable matching for the SMTI and HRT problems, referred to as MAX-SMTI and MAX-HRT, respectively. These problems have been proven to be NP-hard [17, 26]. Equity is another criterion for matchings, especially in societal applications requiring a compromise stable matching [32, 33]. A popular way to assess fairness between the two parties is the sex equality cost [14]. Minimizing this cost for the SM problem, specifically referred to as the sex-equal stable marriage problem (SESM), is proven to be NP-hard [21]. Numerous studies have investigated this problem and proposed corresponding algorithms to address it [16, 20, 37]. However, efficiently solving large instances of MAX-SMTI and MAX-HRT remains an ongoing research topic, with limited studies striving for equity while solving the MAX-SMTI problem.
@@ 1.1 Our Contribution
In this paper, we introduce a tie-breaking based local search algorithm (TBLS) to address the MAX-SMTI and MAX-HRT problems. Initially, TBLS arbitrarily breaks all ties and applies a base algorithm designed for the SM or HR problem, such as the GS algorithm, to achieve a stable matching. The algorithm then iteratively refines the tie-breaking strategy. Rather than adjusting ties randomly, TBLS implements a well-designed adjustment to purposefully introduce a blocking pair (BP) by increasing an agent’s rank within a tie, aimed at filling the unassigned position for this agent. To escape local optima, a disruption that randomly adjusts some ties is applied either when no adjustments are available or at a low probability. Following the tie-breaking strategy refinement process, TBLS uses a BP removal process to effectively secure a stable matching from the existing matching. Instead of re-evaluating the entire agent set, this removal process requires examining only a specific subset of agents to identify and eliminate potential BPs. If this removal process exceeds a predefined time threshold, the base algorithm is reapplied to guarantee a stable matching. This iterative refinement continues until either a perfect matching is found or the maximum number of iterations is reached.
Additionally, we propose an equity-focused variant of TBLS, called TBLS-E, aimed at finding relatively fair matchings for the MAX-SMTI problem. In this variant, an algorithm designed for the SEME problem serves as the base algorithm, and the choice of tie adjustments is strategically restricted to improve fairness while maximizing matching size.
In our experimental study, we refine the random problem generator originally proposed by [13] to generate small-sized and large-sized instances of the MAX-SMTI and MAX-HRT problems. This refined generator produces more challenging instances, as detailed in Section 5.1. We implement ten additional approximation and local search algorithms for comparison. Our experimental results reveal that TBLS consistently outperforms other algorithms in achieving larger matching sizes for both the MAX-SMTI and MAX-HRT problems. Significantly, the equity-focused version, TBLS-E, achieves the lowest sex equality cost for the MAX-SMTI problem while maintaining matching sizes comparable to TBLS. Moreover, both TBLS and TBLS-E exhibit faster computational speed than other local search methods when applied to large-sized problems. Our algorithms introduce a new approach to solving the MAX-SMTI and MAX-HRT problems by iteratively refining tie-breaking strategies. Furthermore, TBLS-E offers a way to obtain a relatively fair matching when solving the MAX-SMTI problem.
@@ 2 Related work
Several polynomial-time approximation algorithms have been proposed to solve the MAX-SMTI and MAX-HRT problems. A straightforward approximation algorithm for these two problems involves applying GS directly after resolving all ties in the given instance. which serves as a 2-approximation algorithm [26]. Building on GS, [22] proposed a $\frac{5}{3}$-approximation algorithm, referred to as Király’s algorithm. Leveraging some ideas from Király’s algorithm, [27] developed the first $\frac{3}{2}$-approximation algorithm, referred to as McDermid’s algorithm. Subsequently, [30] introduced faster and simpler $\frac{3}{2}$-approximation algorithms, referred to as GSM for solving the MAX-SMTI problem and ASBM for solving the MAX-HRT problem. Building on some concepts from GSM and ASBM, [23] proposed natural and local linear-time algorithms, GSA2 and HPA, for solving the MAX-SMTI and MAX-HRT problems, respectively, while maintaining the same approximation ratio. All of the above research focuses on establishing approximation guarantees, without providing any experimental evaluations.
@@
Experimental results are commonly reported in the literature involving constraint programming, local search, or heuristic algorithms. [13] introduced a problem generator for creating random SMTI instances and proposed a constraint programming approach to address the MAX-SMTI problem. [12] developed a local search algorithm called LTIU for the MAX-SMTI problem, which operates by eliminating the undominated blocking pair (UBP). Further refining the local search process, [28] applied the adaptive search (AS)—a meta-heuristic strategy proposed by [6]—to remove UBPs more precisely than relying solely on a single global cost function. For MAX-SMTI problems of size 100, AS is more efficient than both LTIU and McDermid’s algorithm. To efficiently address large-scale MAX-SMTI problems, [38] designed a heuristic search algorithm, MCS, that strategically eliminates UBPs based on maximum conflicts, demonstrating superior performance and lower computational costs compared to AS and LTIU within 3000 iterations. For solving the MAX-HRT problem, the min-conflicts algorithm (MCA) was introduced by [36], which utilizes preference ranks to remove UBPs. Similarly, [3] developed a heuristic repair algorithm (HR), focused on removing UBPs based on preference ranks. Both MCA and HR outperform the LTIU method in effectiveness. [4] proposed a simple heuristic algorithm, called HA, to solve MAX-HRT by gradually forming and removing resident-hospital pairs. Most of the aforementioned literature lacks comparisons with approximation algorithms.
@@
Additionally, extensive research has been done on the SESM problem. A polynomial time approximation algorithm, proposed by [20], seeks to closely approximate the optimal solution. [37] developed a local search algorithm that systematically searches forward from the man-optimal stable matching and backward from the woman-optimal stable matching. [35] innovated the first unbiased and voluntary methods for deriving an equitable stable matching within cubic time, which encompass the EDS, LDS, and PDB algorithms. Few studies have considered optimizing the sex equality cost in the SMTI problem.
@@ 3.1 Bipartite Stable $b$-matching
An undirected bipartite graph $G=(V,E)$, preference lists $L$, and a quota function $b:V\rightarrow\mathbb{N}$ are the inputs of a bipartite stable $b$-matching problem. In the graph $G$, $V=U\cup W$, where $U$ and $W$ are disjoint sets. Each vertex in $U$ is connected to a subset of vertices in $W$, and similarly, vertices in $W$ are connected to subsets in $U$. Each agent $v\in V$ has a preference list $L_{v}$ consisting of agents who are adjacent to $v$. We refer to all the agents in $L_{v}$ as $v$’s candidates. Additionally, the preference list of each agent may include ties. We denote $r_{v}(x)$ as the rank of $x$ in $v$’s preference list. If $v$ prefers $x$ and $y$ equally, this is denoted as $r_{v}(x)=r_{v}(y)$. If $v$ strictly prefer $x$ to $y$, this is denoted as $r_{v}(x)<r_{v}(y)$. Furthermore, define $\phi_{v}(x)=\{y\in L_{x}\mid y\neq v,r_{x}(v)=r_{x}(y)\}$ as a set of agents who have the same rank as $v$ in agent $x$’s preference list.
A $b$-matching $M$ is a subset of $E$ such that each agent $v$ is connected to at most $b(v)$ edges in $M$, that is, $deg_{M}(v)\leq b(v)$, where $deg_{M}$ denotes the degree of vertex $v$ in the graph $G_{M}=(V,M)$. We denote $M(v)$ as the set of agents who are connected to agent $v$ in $M$, referred to as $v$’s partners. An agent $v$ is full if $deg_{M}(v)=b(v)$ and is free if $deg_{M}(v)<b(v)$. The set of all free agents in $M$ is denoted by $F_{M}$. For each free agent $f\in F_{M}$, let $\xi_{f}=\{x\in L_{f}\mid x\notin M(f)\}$ denote the set of agents who are candidates for $f$ but are not currently matched with it. The agent $x$ is termed as $v$’s worst partner if $x\in M(v)$ and $r_{v}(x)=\max\{r_{v}(y)\mid y\in M(v)\}$. In a $b$-matching $M$, $(u,w)$ is a blocking pair (BP) if and only if:
@@ 4 A local search algorithm for SMTI and HRT
In this section, we introduce a tie-breaking based local search algorithm, named TBLS, designed to address the MAX-SMTI and MAX-HRT problems, and its equity-focused variant, TBLS-E. The TBLS algorithm is illustrated in Algorithm 1. TBLS starts to find an efficient strategy $S_{best}$ and a maximum stable matching $M_{best}$ by arbitrarily breaking all the ties and implementing a base algorithm, as outlined in lines 2-5 of Algorithm 1. In each iteration, TBLS performs three steps. First, it refines the current tie-breaking strategy $S$ using the subroutine Refine_Tie_Breaking_Strategy. Following this refinement, it obtains a corresponding stable matching $M$ using the subroutine Obtain_Stable_Matching. Lastly, if $M_{best}$ is found to be worse than $M$ according to the evaluation function $E$, then $M$ and $S$ are assigned to $M_{best}$ and $S_{best}$, respectively, as specified in lines 9-12 of Algorithm 1.
@@
Refine_Tie_Breaking_Strategy refines the current tie-breaking strategy by adjusting the relative order within ties. These adjustments, based on preference ranks and the current stable matching, aim to improve the rank of selected free agents in the preference lists of their respective candidates, thereby increasing their chances of being matched. Each adjustment targets a free agent and one of its candidates, purposefully introducing a BP to alter the current stable matching. When this subroutine is invoked, an adjustment is selected randomly to refine the strategy. Moreover, when no adjustments are available, or with a small probability, the current tie-breaking strategy is disrupted to prevent local optima by randomly adjusting some ties.
Obtain_Stable_Matching obtains a stable matching by either removing BPs or directly using the base algorithm. Given the minor updates to the preference lists during each refinement, attaining a stable matching from the existing matching by removing BPs is typically more efficient than applying the base algorithm. Moreover, this characteristic enables the examination of only a specific subset of agents to identify and eliminate potential BPs, rather than re-evaluating the entire set of agents. Therefore, to effectively obtain a stable matching after the refinement process, a BP removal process is developed to identify and eliminate UBPs. To prevent this process from becoming trapped in an endless cycle, a time threshold is established. Once this threshold is exceeded, the subroutine shifts to applying the base algorithm to secure a stable matching.
The rest of this section is structured as follows: Section 4.1 discusses the evaluation function, Section 4.2 examines the subroutine Refine_Tie_Breaking_Strategy, Section 4.3 explores the subroutine Obtain_Stable_Matching, Section 4.4 outlines two modifications for developing TBLS-E, an equity-focused variant of TBLS.
@@
Input: - An instance $I$. - A small probability $p_{d}$. - A base algorithm $A$. - The maximum number of iterations $max\_iters$. - Two integer $k_{u}$ and $k_{w}$. - The time threshold for obtaining stable matchings $T$.
@@
3 Function Main($I$):
4       $S\mathrel{{:}{=}}$ arbitrarily break all the ties in the instance $I$;
5       $M\mathrel{{:}{=}}$ apply the base algorithm $A$ to find a stable matching after breaking all the ties in the instance $I$ using the tie-breaking strategy $S$;

### v2

@@ Abstract
The stable marriage problem with incomplete lists and ties (SMTI) and the hospitals/residents problem with ties (HRT) are important in matching theory with broad practical applications. In this paper, we introduce a tie-breaking based local search (TBLS) algorithm designed to achieve a weakly stable matching of maximum size for both the SMTI and HRT problems. TBLS begins by arbitrarily resolving all ties and iteratively refines the tie-breaking strategy by adjusting the relative order within ties based on preference ranks and the current stable matching. Additionally, we introduce TBLS-E, an equity-focused variant of TBLS, specifically designed for the SMTI problem. This variant maintains the objective of maximizing matching size, while enhancing equity through two simple modifications. In comparison with ten other approximation and local search algorithms, TBLS achieves the highest matching size, while TBLS-E exhibits the lowest sex equality cost. Significantly, TBLS-E preserves a matching size comparable to that of TBLS. Both our algorithms demonstrate faster computational speed than other local search algorithms in solving large-scale instances. Moreover, our scalability analysis shows that both algorithms maintain efficient performance as problem size increases.
@@ 1 Introduction
A natural variant of SM and HR incorporates ties and incomplete preference lists, allowing the agents to express indifference and exclude unacceptable members. These variants are known as the stable marriage problem with incomplete lists and ties (SMTI) [19] and the hospitals/residents problem with ties (HRT) [17]. Recently, the SMTI and HRT problems have received significant attention from researchers due to their relevance in diverse applications such as the child-adoption matching problem [7, 31] and the student-project allocation problem [1, 2, 25]. When ties are present in the preference lists, three criteria are used to define the stability of a matching: weak stability, strong stability, and super-stability [15, 29]. Among these definitions, weak stability has been the focus of extensive research [4, 7, 13, 11].
Beyond achieving stability, objectives also include maximizing matching size and ensuring fairness. For a given SM or HR problem, all stable matchings have the same size [10]. A stable matching can be easily found using the Gale-Shapley (GS) algorithm proposed by [9]. For the SMTI or HRT problem, a weakly stable matching can be obtained by applying the GS algorithm after breaking all ties. However, different methods of breaking ties generally result in stable matchings of varying sizes [17, 18]. Consequently, a natural objective is to find a maximum cardinality weakly stable matching for the SMTI and HRT problems, referred to as MAX-SMTI and MAX-HRT, respectively. These problems have been proven to be NP-hard [17, 26]. Equity is another criterion for matchings, especially in societal applications requiring a compromise stable matching [32, 33]. A popular way to assess fairness between the two parties is the sex equality cost [14]. The problem of minimizing the sex equality cost in the SM problem is referred to as the sex-equal stable marriage (SESM) problem, which is NP-hard [21]. Numerous studies have investigated this problem and proposed corresponding algorithms to address it [16, 20, 37]. However, efficiently solving large instances of MAX-SMTI and MAX-HRT remains an ongoing research topic, with limited studies striving for equity while solving the MAX-SMTI problem.
@@ 1.1 Our contribution
In this paper, we introduce a tie-breaking based local search (TBLS) algorithm to address the MAX-SMTI and MAX-HRT problems. Initially, TBLS arbitrarily breaks all ties and applies a base algorithm designed for the SM or HR problem, such as the GS algorithm, to achieve a stable matching. Our algorithm then iteratively refines the tie-breaking strategy. Rather than adjusting ties randomly, TBLS implements a well-designed adjustment to purposefully introduce a blocking pair (BP) by increasing an agent’s rank within a tie, aimed at filling the unassigned position for this agent. To escape local optima, a disruption that randomly adjusts some ties is applied either when no adjustments are available or at a low probability. Following the tie-breaking strategy refinement process, TBLS uses a BP removal process to effectively secure a stable matching from the existing matching. Instead of re-evaluating the entire agent set, this removal process requires examining only a specific subset of agents to identify and eliminate potential BPs. If this removal process exceeds a predefined time threshold, the base algorithm is reapplied to guarantee a stable matching. This iterative refinement continues until either a perfect matching is found or the maximum number of iterations is reached.
Additionally, we propose an equity-focused variant of TBLS, called TBLS-E, aimed at finding relatively fair matchings for the MAX-SMTI problem. In this variant, an algorithm designed for the SESM problem serves as the base algorithm, and the choice of tie adjustments is strategically restricted to improve fairness while maximizing matching size.
In our experimental study, we refine the random problem generator originally proposed by [13] to generate small-sized and large-sized instances of the MAX-SMTI and MAX-HRT problems. This refined generator produces more challenging instances by introducing numerous short ties. We implemented ten additional approximation and local search algorithms for comparison. Our experimental results reveal that TBLS consistently outperforms other algorithms in achieving larger matching sizes for both the MAX-SMTI and MAX-HRT problems. Significantly, the equity-focused version, TBLS-E, achieves the lowest sex equality cost for the MAX-SMTI problem while maintaining matching sizes comparable to TBLS. Moreover, both TBLS and TBLS-E exhibit faster computational speed than other local search methods when applied to large-sized problems. Scalability tests also confirm that both algorithms scale efficiently with increasing input sizes. Our algorithms introduce a new approach to solving the MAX-SMTI and MAX-HRT problems by iteratively refining tie-breaking strategies. Furthermore, TBLS-E offers a way to obtain a relatively fair matching when solving the MAX-SMTI problem.
@@ 2 Related work
Several polynomial-time approximation algorithms have been proposed to solve the MAX-SMTI and MAX-HRT problems. A straightforward approximation algorithm for these two problems involves applying the GS algorithm directly after resolving all ties in the given instance. which serves as a 2-approximation algorithm [26]. Building on GS, [22] proposed a $\frac{5}{3}$-approximation algorithm, referred to as Király’s algorithm. Leveraging some ideas from Király’s algorithm, [27] developed the first $\frac{3}{2}$-approximation algorithm, referred to as McDermid’s algorithm. Subsequently, [30] introduced faster and simpler $\frac{3}{2}$-approximation algorithms, referred to as GSM for solving the MAX-SMTI problem and ASBM for solving the MAX-HRT problem. Building on some concepts from GSM and ASBM, [23] proposed natural and local linear-time algorithms, GSA2 and HPA, for solving the MAX-SMTI and MAX-HRT problems, respectively, while maintaining the same approximation ratio. All of the above research focuses on establishing approximation guarantees, without providing any experimental evaluations.
@@
Experimental results are commonly reported in the literature involving constraint programming or heuristic algorithms. [13] introduced a problem generator for creating random SMTI instances and proposed a constraint programming approach to address the MAX-SMTI problem. [12] developed a local search algorithm called LTIU for the MAX-SMTI problem, which operates by eliminating the undominated blocking pair (UBP). Further refining the local search process, [28] applied the adaptive search (AS)—a meta-heuristic strategy proposed by [6]—to remove UBPs more precisely than relying solely on a single global cost function. For MAX-SMTI problems of size 100, AS is more efficient than both LTIU and McDermid’s algorithm. To efficiently address large-scale MAX-SMTI problems, [38] designed a heuristic search algorithm, MCS, that strategically eliminates UBPs based on maximum conflicts, demonstrating superior performance and lower computational costs compared with AS and LTIU within 3,000 iterations. For solving the MAX-HRT problem, the min-conflicts algorithm (MCA) was introduced by [36], which utilizes preference ranks to remove UBPs. Similarly, [3] developed a heuristic repair (HR) algorithm, focused on removing UBPs based on preference ranks. Both MCA and HR outperform the LTIU method in effectiveness. [4] proposed a simple heuristic algorithm, called HA, to solve MAX-HRT by gradually forming and removing resident-hospital pairs. Most of the aforementioned literature lacks comparisons with approximation algorithms in terms of solution quality.
@@
Additionally, extensive research has been done on the SESM problem. A polynomial time approximation algorithm, proposed by [20], seeks to closely approximate the optimal solution. [37] proposed a bidirectional local search algorithm, which simultaneously explores the solution space by conducting a forward search from the man-optimal stable matching and a backward search from the woman-optimal stable matching. At each step, the algorithm employs breakmarriage operations to generate neighboring stable matchings. [35] innovated the first unbiased and voluntary methods for deriving an equitable stable matching within cubic time, which encompass the Late Discontent Suspension (EDS), Early Discontent Suspension (LDS), and Permanent Discontent Ban (PDB) algorithms. Few studies have considered optimizing the sex equality cost in the SMTI problem.
@@ 3.1 Bipartite stable $b$-matching
An undirected bipartite graph $G=(V,E)$, preference lists $L$, and a quota function $b:V\rightarrow\mathbb{N}$ are the inputs of a bipartite stable $b$-matching problem. In the graph $G$, $V=U\cup W$, where $U$ and $W$ are disjoint sets. Each vertex in $U$ is connected to a subset of vertices in $W$, and similarly, vertices in $W$ are connected to subsets in $U$. Each agent $v\in V$ has a preference list $L_{v}$ consisting of agents who are adjacent to $v$. We refer to all the agents in $L_{v}$ as $v$’s candidates. Additionally, the preference list of each agent may include ties. We denote $r_{v}(x)$ as the rank of $x$ in $v$’s preference list. If $v$ prefers $x$ and $y$ equally, this is denoted by $r_{v}(x)=r_{v}(y)$. If $v$ strictly prefer $x$ to $y$, we write $r_{v}(x)<r_{v}(y)$. Furthermore, we define $\phi_{v}(x)=\{y\in L_{x}\mid y\neq v,r_{x}(v)=r_{x}(y)\}$ as a set of agents who have the same rank as $v$ in agent $x$’s preference list.
A $b$-matching $M$ is a subset of $E$ such that each agent $v$ is connected to at most $b(v)$ edges in $M$, that is, $deg_{M}(v)\leq b(v)$, where $deg_{M}$ denotes the degree of vertex $v$ in the graph $G_{M}=(V,M)$. The size of $M$, defined as $|M|$, refers to the number of matched pairs. We denote $M(v)$ as the set of agents who are connected to agent $v$ in $M$, referred to as $v$’s partners. An agent $v$ is full if $deg_{M}(v)=b(v)$ and is free if $deg_{M}(v)<b(v)$. The set of all free agents in $M$ is denoted by $F_{M}$. For each free agent $f\in F_{M}$, let $\xi_{f}^{M}=\{x\in L_{f}\mid x\notin M(f)\}$ denote the set of agents who are candidates for $f$ but are not currently matched with it in matching $M$. A matching $M$ is said to be perfect if and only if there are no free agents.
The agent $x$ is termed as $v$’s worst partner if $x\in M(v)$ and $r_{v}(x)=\max\{r_{v}(y)\mid y\in M(v)\}$. In a $b$-matching $M$, $(u,w)$ is a blocking pair (BP) if and only if:
@@ 4 A local search algorithm for SMTI and HRT
In this section, we introduce a tie-breaking based local search algorithm, named TBLS, designed to address the MAX-SMTI and MAX-HRT problems, and its equity-focused variant, TBLS-E. The TBLS algorithm is illustrated in Algorithm 1. TBLS starts to find an efficient strategy $S_{best}$ and a maximum stable matching $M_{best}$ by arbitrarily breaking all the ties and implementing a base algorithm, as outlined in lines 2-5 of Algorithm 1. In each iteration, TBLS performs three steps. First, it refines the current tie-breaking strategy $S$ using the subroutine RefineTieBreakingStrategy. Following this refinement, it obtains a corresponding stable matching $M$ using the subroutine ObtainStableMatching. Lastly, if $M_{best}$ is found to be worse than $M$ according to the evaluation function $E$, then $M$ and $S$ are assigned to $M_{best}$ and $S_{best}$, respectively, as specified in lines 10-13 of Algorithm 1. TBLS-E is developed by incorporating two key modifications into the original TBLS framework, the details of which are provided in Section 4.4.
In each iteration, the time complexity of TBLS using GS as the base algorithm is $O(n^{2})$, whereas that of TBLS-E using PDB as the base algorithm is $O(n^{3})$. The overall space complexity of both algorithms is $O(n^{3})$. A detailed complexity analysis is provided in Section S1 of Online Resource 1.
@@
RefineTieBreakingStrategy refines the current tie-breaking strategy by adjusting the relative order within ties. These adjustments, based on preference ranks and the current stable matching, aim to improve the rank of selected free agents in the preference lists of their respective candidates, thereby increasing their chances of being matched. Each adjustment targets a free agent and one of its candidates, purposefully introducing a BP to alter the current stable matching. When this subroutine is invoked, an adjustment is selected randomly to refine the strategy. Moreover, when no adjustments are available, or with a small probability, the current tie-breaking strategy is disrupted to prevent local optima by randomly adjusting some ties.
ObtainStableMatching obtains a stable matching by either removing BPs or directly using the base algorithm. Given the minor updates to the preference lists during each refinement, attaining a stable matching from the existing matching by removing BPs is typically more efficient than applying the base algorithm. Moreover, this characteristic enables the examination of only a specific subset of agents to identify and eliminate potential BPs, rather than re-evaluating the entire set of agents. Therefore, to effectively obtain a stable matching after the refinement process, a BP removal process is developed to identify and eliminate UBPs. To prevent this process from becoming trapped in an endless cycle, a time threshold is established. Once this threshold is exceeded, the subroutine shifts to applying the base algorithm to secure a stable matching.
The rest of this section is structured as follows: Section 4.1 discusses the evaluation function, Section 4.2 examines the subroutine RefineTieBreakingStrategy, Section 4.3 presents the subroutine ObtainStableMatching, Section 4.4 outlines two modifications for developing TBLS-E, an equity-focused variant of TBLS.
@@
Input: - An instance II. - A small probability $p_{d}$. - A base algorithm AA. - The maximum number of iterations max_itersmax\_iters. - Two integer $k_{u}$ and $k_{w}$. - The time threshold for obtaining stable matchings $T$.
@@
3 Function Main(II):
4 $S\mathrel{{:}{=}}$ arbitrarily break all the ties in the instance II;
5 $M\mathrel{{:}{=}}$ apply the base algorithm $A$ to find a stable matching after breaking all the ties in the instance II using the tie-breaking strategy $S$;